On December 31, 20x4, you purchase a building for $3091832. The building’s useful life is expected to be 39 years after which it will have a residual value of $552894. The straight-line method of depreciation is used. On December 31, 20x7, the building is sold for $2209078. What is the gain (loss) on sale of the building? Select one: a. $-687451 b. $-441781 c. $32256 d. $-644921
Answer #1

Answer : $- 687,451

Straight line Method of Depreciation

Annual Deprciation = ( cost - residual Value ) / Life in years

Annual Depreciaion = (3091832-552894)/39 =65,100.97

Book value on December 31,20X7 = 3091832 -(65100.97*3) =2896529

Loss on sales of Building = Selling Price - Book value = $ 2209078-2896529. = $- 687,451